OpenStudy (anonymous):

The Atlantic Ocean region contains approximately 2 ×10^16 gallons of water. Lake Ontario has approximately 8,000,000,000 gallons of water. How many Lake Ontarios would it take to fill the Atlantic Ocean region in terms of gallons of water?

OpenStudy (calculusxy):

\[\frac{ 2 \times 10^{16} }{ 8 \times 10^9 }\]

OpenStudy (calculusxy):

notice that i made this into a division expression. it's basically asking how many more times is the atlantic ocean region contains gallons of water compared to lake ontario. do you understand?

OpenStudy (anonymous):

oh I see thank you

OpenStudy (calculusxy):

do you know the answer?

OpenStudy (anonymous):

16

OpenStudy (calculusxy):

\[\frac{ 2 \times 10^{16} }{ 8 \times 10^9 } = \frac{ 2 }{ 8 } \times \frac{ 10^{16} }{ 10^9 } = \frac{ 1 }{ 4 } \times 10^7 = 0.25 \times 10^-1 \times 10^7 = 2.5 \times 10^6\]

OpenStudy (calculusxy):

\[0.25 \times 10^{-1} \times 10^7 = 2.5 \times 10^6\]